OpenStudy (demonchild99):

Find the length of the base of the following pyramid, given the height of the pyramid is 71 meters and the angle of elevation of the base of the pyramid is 42°.

OpenStudy (aaronandyson):

let 1/2 the length = x tan 42 = 71/x x = 78.8 length = 2*78.8 = 157.7 rounded = 158